OpenStudy (anonymous):

1,5,13,25 next in series

OpenStudy (anonymous):

25+16 = 41

OpenStudy (anonymous):

1+4=5 5+8=13 13+12=25 25+16=41

OpenStudy (anonymous):

if u r happy with the answer, pls click GOOD ANSWER button for me :))